OpenStudy (anonymous):

the midpoint of uv is (5 -11) the coordinates of one endpoint are u(3,5) find the coordinates of endpoint v

myininaya (myininaya):

the average of the endpoints is the midpoint

myininaya (myininaya):

that is assume v is (x,y) we have (x+3)/2=5 and (y+5)/2=-11

myininaya (myininaya):

you have two equations to solve

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So all I have to do is find what number plus 3, divided by 2, equals 5? And what number plus 5, divided by two, equals -11? (I'm sorry, I've never been good with math)

myininaya (myininaya):

yep

OpenStudy (anonymous):

Okie dokie!

OpenStudy (anonymous):

x would be 7, right?

OpenStudy (anonymous):

And y would be -22, right?

myininaya (myininaya):

almost (x+3)/2=5 means x+3=10 so x=10-3=7 (y+5)/2=-11 means y+5=-22 so y=-22-5=-27

OpenStudy (anonymous):

Oh, yeah, it would be -27. Thanks! :D

myininaya (myininaya):

np